### v3

@@ 5 Analysis
Figure 6: The coverage ratio comparison between FABind (blue) and FABind+ (pink). Samples are sorted in descending order according to the coverage ratio of FABind+, and the corresponding points are connected to facilitate a clear comparison.
@@
Pocket Radius Module Analysis. As discussed in Section 3.2.1, we introduce dynamic pocket radius prediction to make the pocket cover all atoms of the ligand as much as possible. In this section, we further compare the predicted pockets from FABind and FABind+. We focus on the coverage ratio performance within a subset of the test set, defined by the union of test samples where either FABind or FABind+ could not achieve a 100% coverage ratio. The results, as depicted in Figure 6, reveal that within this collective of test samples, FABind+ demonstrates superior performance, with the coverage ratio for the majority of FABind+ predictions being higher than that of FABind. Such outcomes demonstrate the effectiveness of our proposed dynamic pocket radius prediction. Further analysis is reported in Appendix Section D.
@@
Figure 7: Cases (PDB 6QRG and PDB 6PNN) for permutation-invariant loss analysis. We show the ground truth protein and the predicted conformations by FABind in green and FABind+ in blue.
@@
Figure 8: Case for pocket clustering.
Pocket Clustering Analysis. In this section we conduct a case study to show the effectiveness of our pocket clustering method to detect multiple binding sites, as discussed in Section 3.3.1. We select the protein (PDB 6CKL), which is characterized by its symmetric arrangement of two chains. This symmetry results in two ground truth pockets along with their corresponding ligand conformations. As depicted in Figure 8, with our clustering approach, FABind+ successfully identify both pockets and the corresponding docking conformations through sampling methods. This outcome demonstrates the proficiency of FABind+ in dealing with scenarios involving multiple binding sites, particularly in proteins that form dimers or polymers.
6 Conclusion
In this work, we propose a moleculear docking framework FABind+ based on its previous FABind method. The enhanced FABind+ contains a novel dynamic pocket radius prediction and constrained permutation loss for pose generation. Besides, we also extend the FABind+ to be a sampling version using the dropout trick and a light confidence model. Results show that FABind+ largely outperforms the FABind and achieves superior docking performance. In the future, we will continue exploring better model structures for docking and push the result of alphfafold-latest.
Impact Statements
This paper presents work whose goal is to advance the field of Machine Learning. There are many potential societal consequences of our work, none which we feel must be specifically highlighted here.
@@ Appendix A Experiment Details
A.1 Dataset Processing
We keep the dataset split consistent with previous works Lu et al. (2022); Guo et al. (2023); Pei et al. (2023). The selected 968 validation structures are from data before 2019, and 363 test structures are from 2019 onwards, both ensuring no ligand overlap with the training set. For the training set, we filter out a few samples that could not be read by RDKit or TorchDrug [^4] https://github.com/DeepGraphLearning/torchdrug, leaving 17,795 complexes. We retain protein chains with the nearest atom of the small molecule within 10 Å. Subsequently, samples with protein chains longer than 1500 amino acids, molecules larger than 150 atoms, or inadequate contact (less than 5 amino acids within 10 Å of molecule atoms) were excluded. These criteria resulted in the exclusion of 118, 32, and 1 sample, successively, ultimately yielding 17,644 samples for training.
A.2 Symmetry-Corrected RMSD
In the reported results, the root-mean-square deviation (RMSD) is calculated using sPyRMSD Meli & Biggin (2020). sPyRMSD utilizes a graph matching tool to identify all possible graph isomorphisms, returning the minimum RMSD. This process is aligned with our permutation loss computation. This approach is consistent with our computation of permutation loss. In practice, given the limited number of isomorphic molecules in the test set, we observe a negligible performance gain compared to standard RMSD calculations.
A.3 Permutation Loss Implementation
We utilize the graph-tool toolkit [^5] https://graph-tool.skewed.de/ to identify all symmetric atoms following previous work Zhu et al. (2022). According to sPyRMSD Meli & Biggin (2020), both graph-tool and networkx [^6] https://networkx.org are capable of extracting all permutations. Specifically, an isomorphism between graphs A and B is a bijective mapping of the vertices of graph A to vertices of graph B that preserves the edge structure of the graphs (molecular connectivity in the case of molecular graphs). The problem of finding symmetric atoms can be converted to a graph isomorphism problem.
The implementation of graph-tool is based on VF2 algorithm Cordella et al. (2004). In FABind+, the identification of symmetric atoms is performed offline with multi-processing. This preprocessing takes a few hours to extract all possible permutations of the PDBBind dataset, which is relatively short compared to the training time. Also, it will not increase training cost since the average number of permutations on our used dataset is only $10.1$. The detailed statistics is in Table 3.
Table 3: Statistics on the number of permutations.
@@ A.3 Permutation Loss Implementation
Appendix B Loss Design of Confidence Model
The training objective includes both classification loss, denoted as $L_{cls}$, and ranking loss, denoted as $L_{rank}$. The classification loss is formulated as a cross-entropy loss to accurately predict whether each pose has an RMSD below 2Å. Furthermore, we adopt a pairwise ranking loss inspired by InstructGPT Ouyang et al. (2022). For each batch, samples are copied $N_{copy}$ times, generating $\binom{N_{copy}}{2}$ for each protein-ligand complex. We train on all $\binom{N_{copy}}{2}$ comparisons from each complex as a single batch element. According to InstructGPT, this method is computationally more efficient and reduces the likelihood of overfitting. The ranking loss is formally defined as follows:
$L_{rank}=-\frac{1}{\binom{N_{copy}}{2}}E_{\left(R_{+},R_{-}\right)\sim D}\left[\log\left(\sigma\left(f_{\theta}\left(R_{+}\right)-f_{\theta}\left(R_{-}\right)\right)\right)\right],$ (3)
where $f_{\theta}(R)$ is the scalar score predicted by the confidence model for conformation $R$, $R_{+}$ is the preferred conformation with lower RMSD compared to $R_{-}$, and $D$ is the dataset of comparison sets.
In the implementation, each GPU batch consists of $N_{sample}=1$ samples, which are then duplicated $N_{copy}=8$ times. Following model forward propagation, each sample yields $N_{copy}$ predicted structures. These structures are subsequently ranked according to their computed RMSD values, which are also utilized to calculate the classification loss $L_{cls}$. The total loss for the confidence model, denoted as $L_{conf}$, is the sum of the classification loss and the ranking loss: $L_{conf}=L_{cls}+L_{rank}$.
@@ Appendix C Training and Inference
MLP Configuration. All multilayer perceptrons (MLPs) implemented in FABind+ consist of a Layer Normalization, followed by two linear transformations with ReLU activations. An additional ReLU activation is applied after the final linear transformation if the output of the MLP is an embedding. To regulate the parameters of FABind+, the MLP hidden scale is set to 1, indicating that the MLP’s hidden size is the same as the input embedding dimension. On the other hand, in the confidence model, the MLP hidden scale is increased to 4 to maximize the model capacity.
Training Details for FABind+. Our FABind+ models are trained on eight NVIDIA Tesla V100 GPUs for 1500 epochs. We use Adam Kingma & Ba (2014) as the optimizer, and set the hyperparameter $\epsilon$ to $81e-$ 8 and $(\beta_{1},\beta_{2})$ to $(0.9,0.999)$ with no weight decay. The peak learning rate is set to $55e-$ 5 with a 15-epoch warmup stage followed by a linear decay learning rate scheduler. The dropout probability and total batch size are set to $0.1$ and $16$.
Table 4: The hyperparameter options we searched through for FABind+. The final parameters are marked in bold.
@@
Table 5: The hyperparameter options we searched through for confidence model. The final parameters are marked in bold.
@@
model backbone MLP, stacked MLP, node MLP, FABind layer
@@
Training Details for Confidence Model. The computational overhead for confidence model training is notably lower, requiring 15 epochs of training on eight NVIDIA Tesla V100 GPUs. The detailed hyperparameter options are listed in Table 5. We evaluate various model backbones for feature extraction, namely MLP, Stacked MLP, Node MLP, and FABind layer. Stacked MLP refers to the configuration where two previously described MLPs are stacked, thus expanding two sub-layers to four. Node MLP denotes an additional MLP transformation applied to node embeddings prior to their aggregation (sum). Subsequent to aggregation, another MLP updates the features for the output. Incorporating an additional FABind layer results in the highest parameter count. However, we observe limited benefit in enhancing the confidence model’s capability. This underscores that generalization is crucial for the efficacy of the confidence model.
@@
During training, our confidence model can generate training examples on-the-fly, in contrast to the approach in DiffDock Corso et al. (2023b), which requires offline storage of sampled structures. This difference arises because DiffDock treats the scoring and confidence models as separate models, whereas our method integrates them into a single model. With the sampling mode on, the confidence model can directly utilize samples generated by FABind+. Consequently, an MLP-based backbone suffices due to the robust feature extraction capabilities of the preceding FABind+.
Appendix D Further Analysis
D.1 Pocket Prediction Study
Figure 9: Analysis of pocket prediction and the predicted ligand RMSD. “Max Distance” is the distance between the predicted pocket center and the farthest ground truth ligand atom.
D.2 Ablation Study
Figure 10: Ablation of model size.
Ablations for Regression FABind+. To transition from FABind to a regression-based method, we implemented several enhancements depicted in Figure 14. The introduction of radius prediction and permutation loss contributed to performance increases of 3.0% and 2.1%, respectively. A thorough hyperparameter search further unlocked the model’s capabilities, yielding an additional 2.1% improvement. Given the presence of six loss functions in our framework, fine-tuning the weights assigned to each loss function had a significant impact on the results. Moreover, we discovered that simply increasing the number of layers could also lead to a performance gain. However, it is important to note that deeper architectures require a more careful parameter study.
Figure 11: Performance gain from FABind to FABind+.
Ablations for Sampling FABind+. For the sampling version, we present a scaling curve in Figure 12 to illustrates how performance changes for different ablations. Here, “FABind+” represents our sampling-based model. “No Dropout” refers to our model operating without dropout sampling during both training and inference phases, whereas “No Clust” indicates the absence of pocket clustering.
This figure underscores the importance of each design element in achieving optimal performance. Specifically, removing pocket clustering leads to a slight drop in performance, as it hinders the ability to capture multiple binding sites. The lack of dropout sampling means that diversity relies entirely on pocket clustering. However, given the limited number of pockets, the orange dashed line in the graph exhibits only a minimal increase with sample size.
Figure 12: Ablation analysis for the sampling version.
Appendix E Visualizations
@@
Figure 13: Case studies for Regression-based FABind+. Structures predicted by FABind+ (green), FABind (magentas), DiffDock (blue), TankBind (yellow) and EquiBind (orange) are placed together with the protein target, with the RMSD to the ground truth (red) reported. These comparisons underscore the capability of FABind+ for accurate prediction.
Figure 14: Sampling diversity and quality with FABind+. On the left, sampled structures (green) alongside the ground truth (red) are depicted; on the right, a closer view of the sampled structures is provided, each colored randomly for distinction. (a) In the case of complex 5ZCU, which features a single binding site, FABind+ successfully generates a variety of conformations that are centered around the ground truth. (b) For complex 6AGT, characterized by multiple pockets, FABind+ not only identifies these pockets but also generates diverse structures within each pocket.

### v4

@@ 5.1. Component Analysis
Figure 6. Coverage ratio comparison between FABind (blue) and FABind+ (pink). Samples are sorted in descending order according to the coverage ratio of FABind+, with points connected for clarity.
@@
Pocket Radius Module Analysis. As discussed in Section 3, we introduce dynamic pocket radius prediction to make the pocket cover all atoms of the ligand as much as possible. Here we further compare the predicted pockets from FABind and FABind+. We focus on the coverage ratio performance within a subset of the test set, defined by the union of test samples where either FABind or FABind+ could not achieve a 100% coverage ratio. The results, as depicted in Figure 6, reveal that within this collective of test samples, FABind+ demonstrates superior performance, with the coverage ratio for the majority of FABind+ predictions being higher than that of FABind. Such outcomes demonstrate the effectiveness of our proposed dynamic pocket radius prediction.
Figure 7. Analysis of pocket prediction and the predicted ligand RMSD in the test set. “Max Distance” is the distance between the predicted pocket center and the farthest ground truth ligand atom.
@@
Figure 8. Cases (PDB 6QRG and PDB 6PNN) for permutation-invariant loss analysis. We show the ground truth protein and the predicted conformations by FABind in green and FABind+ in blue.
@@
Figure 9. Case for pocket clustering.
Pocket Clustering Analysis. Here we conduct a case study to show the effectiveness of our pocket clustering method in detecting multiple binding sites, as discussed in Section 3.3.1. We select the protein (PDB 6CKL), which is characterized by its symmetric arrangement of two chains. This symmetry results in two ground truth pockets along with their corresponding ligand conformations. As depicted in Figure 9, with our clustering approach, FABind+ successfully identifies both pockets and the corresponding docking conformations through sampling methods. In contrast, regression models are unable to capture this scenario. This case underscores the proficiency of the sampling-based FABind+.
Visualizations of Sampling Capability. We also showcasing the sampling results in Figure 10. In the case of complex 5ZCU, which features a single binding site, FABind+ successfully generates a variety of conformations that are centered around the ground truth. For complex 6AGT, characterized by multiple pockets, we can observe that FABind+ not only identifies these pockets but also generates diverse structures within each pocket.
Figure 10. Sampling diversity of FABind+. On the left, sampled structures (green) alongside the ground truth (red) are depicted; on the right, a closer view of the sampled structures is provided, each colored randomly for distinction.
5.2. Ablation Study
Table 3. Ablation study on ligand RMSD metric.
[…]
regression FABind+ 43.5% 5.2 2.6
- w/o. dynamic rad 40.0% 5.5 2.8
- w/o. permutation loss 41.2% 5.6 2.8
- w/o. add one layer 39.1% 5.5 2.7
sampling FABind+ (10) 42.4% 5.2 2.6
- w/o. confidence model 35.3% 6.4 3.1
- classification loss only 41.9% 5.4 2.6
- ranking loss only 41.7% 5.5 2.7
We conducted a detailed ablation study to evaluate the impact of various components on both the regression and sampling versions of FABind+. The results are summarized in the Table 3.
For the regression version, removing the dynamic radius adjustment (‘- w/o. dynamic rad‘), permutation loss (‘- w/o. permutation loss‘), or additional layer (‘- w/o. add one layer‘) resulted in accuracy drops, with the success rate for predictions under 2Å decreasing from 43.5% to 40.0%, 41.2%, and 39.1%, respectively. The mean and median RMSD also worsened without these design elements.
In the sampling version (with a sample size of 10), the absence of the confidence model (‘- w/o. confidence mode‘) caused a significant drop in performance, reducing the success rate from 42.4% to 35.3%. We also compared the effects of using only classification loss or only ranking loss. The results showed that while both losses contribute to performance, using them together yields the best results.
Figure 11. Ablation analysis for the sampling version.
Ablations for Sampling Capability. For the sampling version, we additionally present a scaling curve in Figure 11 to illustrates how performance changes for different ablations. This figure underscores the importance of each design element in achieving optimal performance. Specifically, removing pocket clustering (‘No Clust‘) leads to a slight drop in performance, as it hinders the ability to capture multiple binding sites. The lack of dropout sampling (‘No Dropout‘) means that diversity relies entirely on pocket clustering. However, given the limited number of pockets, the orange dashed line in the graph exhibits only a minimal increase with sample size.
6. Conclusion
In this work, we demonstrate that FABind+ can serve as a unified model for both regression and sampling. We enhance the original FABind method by introducing dynamic pocket radius prediction and permutation loss. Additionally, we transform the pre-trained regression model into a sampling model using pocket clustering and dropout sampling without further training. This approach to activating sampling capability is universal to all regression-based models. A lightweight confidence model is then trained to select the best conformation. Our results show that FABind+ significantly outperforms FABind, achieving superior docking performance.
Acknowledgements.
This work is supported by National Natural Science Foundation (62076105).
@@ Appendix A Experiment Details
A.1. Dataset Processing
We keep the dataset split consistent with previous works (Lu et al., 2022; Guo et al., 2023; Pei et al., 2023). The selected 968 validation structures are from data before 2019, and 363 test structures are from 2019 onwards, both ensuring no ligand overlap with the training set. For the training set, we filter out a few samples that could not be read by RDKit or TorchDrug [^4] https://github.com/DeepGraphLearning/torchdrug, leaving 17,795 complexes. We retain protein chains with the nearest atom of the small molecule within 10 Å. Subsequently, samples with protein chains longer than 1500 amino acids, molecules larger than 150 atoms, or inadequate contact (less than 5 amino acids within 10 Å of molecule atoms) were excluded. These criteria resulted in the exclusion of 118, 32, and 1 sample, successively, ultimately yielding 17,644 samples for training.
A.2. Symmetry-Corrected RMSD
In the reported results, the root-mean-square deviation (RMSD) is calculated using sPyRMSD (Meli and Biggin, 2020). sPyRMSD utilizes a graph matching tool to identify all possible graph isomorphisms, returning the minimum RMSD. This process is aligned with our permutation loss computation. This approach is consistent with our computation of permutation loss. In practice, given the limited number of isomorphic molecules in the test set, we observe a negligible performance gain compared to standard RMSD calculations.
A.3. Permutation Loss Implementation
Table 4. Statistics on the number of permutations.
@@ A.3. Permutation Loss Implementation
We utilize the graph-tool toolkit [^5] https://graph-tool.skewed.de/ to identify all symmetric atoms following previous work (Zhu et al., 2022). According to sPyRMSD (Meli and Biggin, 2020), both graph-tool and networkx [^6] https://networkx.org are capable of extracting all permutations. Specifically, an isomorphism between graphs A and B is a bijective mapping of the vertices of graph A to vertices of graph B that preserves the edge structure of the graphs (molecular connectivity in the case of molecular graphs). The problem of finding symmetric atoms can be converted to a graph isomorphism problem.
The implementation of graph-tool is based on VF2 algorithm (Cordella et al., 2004). In FABind+, the identification of symmetric atoms is performed offline with multi-processing. This preprocessing takes a few hours to extract all possible permutations of the PDBBind dataset, which is relatively short compared to the training time. Also, it will not increase training cost since the average number of permutations on our used dataset is only $10.1$. The detailed statistics are in Table 4.
Appendix B Model Details of FABind+
MLP Configuration. All multilayer perceptrons (MLPs) implemented in FABind+ consist of a Layer Normalization, followed by two linear transformations with ReLU activations. An additional ReLU activation is applied after the final linear transformation if the output of the MLP is an embedding. To regulate the parameters of FABind+, the MLP hidden scale is set to 1, indicating that the MLP’s hidden size is the same as the input embedding dimension. On the other hand, in the confidence model, the MLP hidden scale is increased to 4 to maximize the model capacity.MLP Configuration. All multilayer perceptrons (MLPs) implemented in FABind+ consist of a Layer Normalization, followed by two linear transformations with ReLU activations. An additional ReLU activation is applied after the final linear transformation if the output of the MLP is an embedding. To regulate the parameters of FABind+, the MLP hidden scale is set to 1, indicating that the MLP’s hidden size is the same as the input embedding dimension. On the other hand, in the confidence model, the MLP hidden scale is increased to 4 to maximize the model capacity.
Figure 12. Exploration of model size.
Implementation Details. Initial ligand conformations are generated using the ETKDG algorithm (Riniker and Landrum, 2015), followed by MMFF optimization via RDKit (Landrum et al., 2013), which produces the random generation of a low-energy ligand conformation. For pocket radius prediction, considering the fixed radius of FABind is 20Å, we adjust any predicted radius falling below this 20Å threshold to 20Å. This adjustment ensures the predicted pocket is at least as large as those predicted by FABind. For other hyperparameters and training configurations, we put the details in Appendix C.
@@ Appendix C Training and Inference
Table 5. The hyperparameter options we searched through for FABind+. The final parameters are marked in bold.
@@
Training Details for FABind+. Our FABind+ models are trained on eight NVIDIA Tesla V100 GPUs for 1500 epochs. We use Adam (Kingma and Ba, 2014) as the optimizer, and set the hyperparameter $\epsilon$ to $81e-$ 8 and $(\beta_{1},\beta_{2})$ to $(0.9,0.999)$ with no weight decay. The peak learning rate is set to $55e-$ 5 with a 15-epoch warmup stage followed by a linear decay learning rate scheduler. The dropout probability and total batch size are set to $0.1$ and $16$.
Table 6. The hyperparameter options we searched through for confidence model. The final parameters are marked in bold.
@@
model backbone MLP, stacked MLP, node MLP, node Attention, FABind layer
@@
Training Details for Confidence Model. The computational overhead for confidence model training is notably lower, requiring 15 epochs of training on eight NVIDIA Tesla V100 GPUs. The detailed hyperparameter options are listed in Table 6. We evaluate various model backbones for feature extraction, namely MLP, Stacked MLP, Node MLP, Node Attention and FABind layer. Stacked MLP refers to the configuration where two previously described MLPs are stacked, thus expanding two sub-layers to four. Node MLP denotes an additional MLP transformation applied to node embeddings prior to their aggregation (sum). Subsequent to aggregation, another MLP updates the features for the output. Node Attention is similar to Node MLP but includes a self-attention module before aggregation. FABind layer results in the highest parameter count. However, we observe limited benefits in enhancing the confidence model’s capability. This underscores that generalization is crucial for the efficacy of the confidence model.
@@
During training, our confidence model can generate training examples on-the-fly, in contrast to the approach in DiffDock (Corso et al., 2023b), which requires offline storage of sampled structures. This difference arises because DiffDock treats the scoring and confidence models as separate models, whereas our method integrates them into a single model. With the sampling mode on, the confidence model can directly utilize samples generated by FABind+. Consequently, an MLP-based backbone suffices due to the robust feature extraction capabilities of the preceding FABind+.
Appendix D More Visualizations
@@
Figure 13. Case studies for Regression-based FABind+. Structures predicted by FABind+ (green), FABind (magentas), DiffDock (blue), TankBind (yellow) and EquiBind (orange) are placed together with the protein target, with the RMSD to the ground truth (red) reported. These comparisons underscore the capability of FABind+ for accurate prediction.